OpenStudy (anonymous):

Samantha swam upstream for some distance in one hour. She then swam downstream the same river for the same distance in only 12 minutes. If the river flows at 4 mph, how fast can Samantha swim in still water?

OpenStudy (anonymous):

@moeshad12 \(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e~\color{#00ff00}t\color{#00ff49}o~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)

OpenStudy (anonymous):

Do you know the formula for Upstream,Downstream?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

r t = d Let s be Samantha's speed in mph in still water. Solve the following for s.\[1 (s-4)=\frac{s+4}{5} \] x = 6

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Sorry, x = 6 should have been submitted as s = 6.